help

#5 a-d


5. (3 pts each) Let (x) = 3 - 2x and g(x) = x². Evaluate each function. a. f(-4) b. ga) c. g(a +3) d. f(2) - 9(-3)
Answer #1

Heroes fca) = 3-24 and JE 4) = 3 – 2 (-4) = 3 +8 g(x) = x² HE 4) = 11 gea) (a)? = Q² gea+3) = (+3) a +60 +9 8 (2) 94-3). (-3)